# Recipes for Pre-training LLMs with MXFP8

### v1

@@ 4 Related work
Low-precision training and inference is a widely studied topic in deep learning literature. While significant progress has been made in low-precision inference [22, 23, 24, 25], there are relatively fewer studies demonstrating low-precision techniques for pre-training LLMs, especially large-scale pre-training on large token horizon. Our work primarily focuses on 8-bit pre-training and prior work on related low-precision LLM pre-training can be grouped into the following two categories:
  1. LLM pre-training using FP8 formats: [6] proposes an FP8 binary interchange format, consisting of E4M3 and E5M2 encodings, and a per-tensor scaling approach — an entire tensor is scaled to capture the maximum number of tensor values in the representable range in FP8[26] discusses FP8 pre-training challenges and proposes model-level modifications to train a 7B parameter model. Recently, per-tensor scaled FP8 was used to train the Nemotron-H family of LLMs [8] and the Llama-4 family of models also used FP8  [27]. Instead of per-tensor scaling, DeepSeek-V3 family of models [10] use block-scaled FP8, this helps to better capture outliers and minimize quantization errors.
  2. Pre-training using MXFP formats: [2] presents empirical data on pre-training models with MXFP formats. They show cast-only inference results for MXFP8 and pre-training results for MXFP6 and MXFP4-weights. [28] studies MXFP4 backward-pass quantization and [29] investigates MXFP4 weight quantization on relatively small token horizon. All these studies are based on the scaling factor computation method described in [1] which we show to be ineffective at large token horizons.
@@ 5 Conclusions and future work
As part of our future work, we plan to extend our study to post-training phases (e.g. phases that perform RL based policy-optimizations and its variants [31] as well as supervised-finetuning stages). Additionally, from a hardware systems performance perspective, since tensors have to be quantized along two separate axes with MX-formats, the training framework needs to store two copies for each tensor (we discussed this in section 3.2). We plan to study scaling methods (e.g. 2D scaling instead of 1D) that would lower the storage overhead.
@@ Acknowledgments
We thank members of the ADLR/PSX team (Sweta P., Mikail K., Ben L.) for helping with the draft revisions as well as Mohammad S., Carlo D.M., Michael A., Eric C. and Bryan C. with valuable feedback and discussions. We also thank PM for guidance throughout this work.
@@ A.1 UE8M0 rounding
Computation described in Algorithm 1 is a simplification. The simplification comes from storing the output of $log_{2}$(float x) in FP32. $log_{2}$ function in device internally performs a round-to-nearest of the resulting return value and thus the result of ceil($log_{2}$(x)) can be different if the output of $log_{2}$ is not stored in a sufficiently large data-type. Hence, for emulation purposes we work directly with the bit representation of the ratio of amax and destmax. We, next, describe the computation flow.
Background: As a reminder, the quantization process from 32 high-precision values, $V_{i}$, to quantized values, $Q_{i}$; $1\leq i\leq 32$, is given by: $Q_{i}$ = Quantize_to_fp8$(V_{i}/2^{X})$. $2^{X}$ is the scale factor; $X$ is stored in an unsigned 8bit integer container in memory and interpreted as $2^{X}$ by the hardware. This scale factor decodes $Q_{i}$ back to $V_{i}$ (with quantization loss).
Value of scale factor ($2^{X}$) = float_to_8bits(amax/destmax), where amax is absolute maximum in input/source block (of 32 elements) and destmax is the largest positive number in the destination (MX) number system. float_to_8bits converts a floating-point number to a power-of-two number.
A float (FP32) number can be represented in IEEE convention as $2^{E}$$\times$ 1.mantissa (normal) or $2^{-126}$$\times$ 0.mantissa (denormal). E can lie between -127 to 127 (or 0 to 254 with the exponent bias) and can be represented in the 8-bit container for scale factor. -126 (or 1 with the exponent bias) is also representable by 8-bit container. mantissa lies between [0,1). So, the question is: should mantissa bits be rounded-up, rounded-down, round-to-nearest, discarded, etc. to create a power-of-two number? We find round-up to be the best choice for pre-training with MX-formats.
@@ A.2 MX-format conversion
Section 2 relies on the standard MX-format conversion algorithm defined in [2], but for completeness we show it here given a shared scaling exponent $X$ as computed in 3.1.
Quantizing FP32 values to MX type: Once X is computed, $V_{i}/2^{X}$ is computed and the resulting value is quantized to a FP8-representable number (Quantize_to_fp8()). Round-to-nearest-ties-to-even (RN) rounding is used during this quantization step. The conversion process is saturating, i.e. if after rounding the resulting value exceeds FP8 max or is less than FP8 min value, then the result is clamped to respective max or min value.
@@ A.4 MXFP8 pre-training for a mixture-of-experts model
Section 3.3 discusses empirical data that shows MXFP8 matches BF16 accuracy (both training loss as well as downstream task accuracy). Transformer based mixture-of-experts (MoE) models are popular in literature. Fig. 6 shows that MXFP8 pre-training also matches BF16 pre-training loss curve for a MoE setup that we experimented with. The MoE model has 16 billion total parameters and $\sim$2.5 billion active parameters and we train the model on 1 trillion tokens. We follow the same guidelines discussed in section 3.2 for pre-training the MoE model. The pre-training phase uses a WSD [32] learning rate schedule. The final loss of the MXFP8-trained MoE model is within 0.1% of BF16-training.

### v2

@@ 4 Related work
Low-precision training and inference is a widely studied topic in deep learning literature. While significant progress has been made in low-precision inference [24, 25, 26, 27], there are relatively fewer studies demonstrating low-precision techniques for pre-training LLMs, especially large-scale pre-training on a large token horizon. Our work primarily focuses on 8-bit pre-training and prior work on related low-precision LLM pre-training can be grouped into the following two categories:
  1. LLM pre-training using FP8 formats: [7] proposes an FP8 binary interchange format, consisting of E4M3 and E5M2 encodings, and a per-tensor scaling approach — an entire tensor is scaled to capture the maximum number of tensor values in the representable range in FP8[28] discusses FP8 pre-training challenges and proposes model-level modifications to train a 7B parameter model. Recently, per-tensor scaled FP8 was used to train the Nemotron-H family of LLMs [9] and the Llama-4 family of models also used FP8  [29]. Both of these approaches mention the need to keep a few transformer blocks in high-precision. Instead of per-tensor scaling, DeepSeek-V3 family of models [22] use block-scaled FP8, this helps to better capture outliers and minimize quantization errors. With block-scaling setup, certain tensors require vector 1x128 scaling and certain tensors require per-block (e.g. 128x128) software scaling. Native support for MXFP8 simplifies this — fine-grained scaling provides better numerical robustness and hardware support for scaling avoids any tradeoff between smaller block sizes and hardware speed.
  2. Pre-training using MXFP formats: [3] presents empirical data on pre-training models with MXFP formats. They show cast-only inference results for MXFP8 and pre-training results for MXFP6 and MXFP4-weights. [30] studies MXFP4 backward-pass quantization and [31] investigates MXFP4 weight quantization on relatively small token horizon.
@@ 5 Conclusions and future work
As part of our future work, we plan to extend our study to post-training phases (e.g. phases that perform reinforcement learning (RL)-based policy optimizations and their variants [33] as well as supervised fine-tuning stages). Additionally, from a hardware systems performance perspective, since tensors have to be quantized along two separate axes with MX-formats, the training framework needs to store two copies for each tensor (we discussed this in section 3.2). We plan to study scaling methods that would lower the storage overhead.
@@ Acknowledgments
We thank Sweta Priyadarshi, Mikail Khona and Ben Lanir for helping with the draft revisions as well as Mohammad Shoeybi, Carlo del Mundo, Michael Andersch, Eric Chung and Bryan Catanzaro for valuable feedback and discussions.
@@ A.1 UE8M0 rounding
Computation described in Algorithm 1 is a simplification. The simplification comes from storing the output of log2(float x) in FP32. log2 function on device internally performs a round-to-nearest of the resulting return value and thus the result of ceil(log2(x)) can be different if the output of log2 is not stored in a sufficiently large data-type. Hence, for emulation purposes we work directly with the bit representation of the ratio of amax and destmax. We, next, describe the computation flow.
[…]
Value of scale factor ($X$) = float_to_8bits(amax/destmax), where amax is absolute maximum in input/source block (of 32 elements) and destmax is the largest positive number in the destination (MX) number system. float_to_8bits converts a floating-point number to a power-of-two number.
A float (FP32) number can be represented in IEEE convention as $2^{E}$ $\times$ 1.mantissa (normal) or $2^{-126}$ $\times$ 0.mantissa (subnormal). E can lie between -127 to 127 (or 0 to 254 with the exponent bias) and can be represented in the 8-bit container for scale factor. -126 (or 1 with the exponent bias) is also representable by 8-bit container. mantissa lies between [0,1). So, the question is: should mantissa bits be rounded-up, rounded-down, round-to-nearest, discarded, etc. to create a power-of-two number? We find round-up to be the best choice for pre-training with MX-formats.
@@ A.2 MX-format conversion
Section 2 relies on the standard MX-format conversion algorithm defined in [3], but for completeness we show it here given a shared scaling exponent $X$ as computed in 3.3.
Quantizing FP32 values to MX type: After X is computed, $V_{i}/X$ is computed and the resulting value is quantized to a FP8-representable number (Quantize_to_fp8()). Round-to-nearest-ties-to-even (RN) rounding is used during this quantization step. The conversion process is saturating, i.e. if after rounding the resulting value exceeds FP8 max or is less than FP8 min value, then the result is clamped to respective max or min value.
@@ A.4 MXFP8 pre-training for a mixture-of-experts model
Section 3.1 discusses empirical data that shows MXFP8 matches BF16 accuracy (both training loss as well as downstream task accuracy). Transformer based mixture-of-experts (MoE) models are popular in literature. Figure 6 shows that MXFP8 pre-training also matches BF16 pre-training loss curve for a MoE setup that we experimented with. The MoE model has 16 billion total parameters and $\sim$2.5 billion active parameters and we train the model on 1 trillion tokens. We follow the same guidelines discussed in section 3.2 for pre-training the MoE model. The pre-training phase uses a WSD [34] learning rate schedule. The final loss of the MXFP8-trained MoE model is within 0.1% of BF16-training.